OpenStudy (nirmalnema):

why 0! and 1! both are equal to one

terenzreignz (terenzreignz):

Because we have this identity :) \[\Large n! = n\cdot (n-1)!\] So, by that logic, we have: \[\Large 1! = 1(1-1)!\] \[\Large 1 = 1\cdot0!\]

terenzreignz (terenzreignz):

Implying that 0! = 1

OpenStudy (nirmalnema):

ok.... BUT HOW 0! is = to 1

terenzreignz (terenzreignz):

Okay... THIS \[\Large n!= n(n-1)!\] is true, right? Just take n = 1.

OpenStudy (anonymous):

You pretty much just have to accept it. It simplifies very complex concepts

terenzreignz (terenzreignz):

\[\Large 1! = 1\cdot(1-1)!\] Now, we know that 1! = 1 \[\Large 1 = 1\cdot 0!\] Now, solving for 0! does yield \[\Large 0! = 1\]

OpenStudy (nirmalnema):

ok..thank you